Calculate the volume, in milliliters, of 0.170 M NaOH that must be added to 283 mL 0.0509 M 3‑(?‑Morpholino) propanesulfonic acid (MOPS) to give the solution a pH of 7.55. The p?a of MOPS is 7.18.
Answer #1

let the volume of NaOH = v in ml

no of moles of NaOH   = molarity * volume in ml

                                     = 0.170*v

no of moles of MOPS = molarity * volume in ml

                                     = 0.0509*283 = 14.4047 mmoles

   MOPS (aq) + NaOH(aq) ----------> NaMOPS + H2O(l)

I         14.4047             0.170v                      0

C       -0.170v          -0.170v                    0.170v

E       14.4047-0.170v     0                          0.170v

      PH   = PKa + log[NaMOPS]/[MOPS]

        7.55   = 7.18 + log0.170v/(14.4047-0.170v)

     7.55-7.18    = log0.170v/(14.4047-0.170v)

   0.37   = log0.170v/(14.4047-0.170v)

log0.170v/(14.4047-0.170v)   = 0.37

0.170v/(14.4047-0.170v)    = 2.3442

0.170v   = 2.3442*(14.4047-0.170v)

   v   = 59.4ml

The volume of NaOH = 59.4ml   >>>answer